OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x - 3y = -11 2x - 6y = -14 -1 zero 1 2

OpenStudy (anonymous):

(2) 5x-3y=-11 (-5)2x-6y=-14 10x-6y=-22 -10x+30y=70 24y=48 y=48/24 y=2